When solving for x in the equation


((9x + 7)/2)-(x-(x-2))/7)=36



the equation that is NOT correct, is



[1] ((9x + 7)/2)-[(7x/7)-((x-2)/7)]=36



[2] ((9x + 7)/2)-((7x-x-2)/7)=36



[3] ((9x + 7)/2)-((6x+2)/7)=36



[4] 7(9x + 7) − 2(6x + 2) = 14 × 36.



1
Expert's answer
2022-04-17T23:11:10-0400

We have the equation

((9x + 7)/2)-(x-(x-2)/7)=36


When solving for x there will be next steps:

Step1. Bring "x - (x-2)/7" to a common denominator 7.

((9x + 7)/2)-[(7x/7)-((x-2)/7)]=36


Step 2. Write "(7x/7)-((x-2)/7)" as one fraction.

((9x + 7)/2)-((7x-x+2)/7)=36


Step 3. Simplify the numerator of the fraction "(7x-x+2)/7".

((9x + 7)/2)-((6x+2)/7)=36


Step 4. Multiply the left and right parts of the equation by the lcm of its denominators.

7(9x + 7) − 2(6x + 2) = 14 × 36.


Comparing our steps with the variants in the task we can see that in [2] there is one wrong sign ("-" before 2 when it should be "+").


Answer: The equation that is NOT correct, is [2].